Please explain to me how i am supposed to graph this equation: 1.5x+y =4
Answer by stanbon(75887)

Draw an x-y axis system.
Now look at your equation:
Let x=0, then y =4
That gives you a point (0,4).
Now let y=0, then x=4/1.5 = 8/3
That gives you a point (8/3,0)
Now plot the two points on your x-y axis system.
Now draw a line through the two points.
Cheers,
Stan H.